# BillBoard Splatting (BBSplat): Learnable Textured Primitives for Novel View Synthesis

### v3

@@ 4.3 Ablation study
In Table 5, we provide an ablation study of different aspects of our approach. Namely the use of RGB and alpha textures, texture regularization, and compression post-processing.
The use of both alpha and RGB textures is essential for rendering quality. Excluding $T^{\alpha}$ significantly increases LPIPS, while excluding $T^{\textrm{RGB}}$ reduces PSNR and SSIM. Texture regularization term preserves over-fitting and leads to additional metrics improvement. It also encourages sparsity in textures and reduces storage costs in combination with compression. The use of compression leads to a slight metric reduction but allows to significantly reduce storage space requirements. Therefore, all parts of the proposed approach are essential to get the best result. However, we note that metrics could be additionally improved by disabling compression.
@@ 5 Conclusion
Limitations and future work. One of the main limitations of the proposed method is storage space. While we make a significant step towards its reduction, there is still room for improvement in future works. Another limitation is training time, as it takes about 40 minutes to fit a scene, compared with 5 minutes for 3DGS. This slowdown is caused by backpropagation to the textures, and while it still significantly outperforms NeRF-based methods in terms of speed, it could be a bottleneck for some applications. In future works, we are going to focus on resolving these limitations.
@@ 7.2 Detailed results
In Table 8, we report metrics for all scenes. For competitors, we provide metrics for the maximum recommended number of Gaussians, and for ours with 300K billboards. The averaged values for these three datasets are reported in the main paper. For examples of renderings from different angles, see the video in the supp. mat..
@@
Tanks&Temples Train 22.57 / 0.8248 / 0.1968 23.26 / 0.8531 / 0.1735 22.03 / 0.8146 / 0.2144 23.04 / 0.8336 / 0.1696
Truck 26.04 / 0.8851 / 0.1489 27.23 / 0.9024 / 0.1230 26.07 / 0.8814 / 0.1612 26.06 / 0.8800 / 0.1336
Francis 28.00 / 0.9127 / 0.2349 29.24 / 0.9217 / 0.2275 28.23 / 0.9118 / 0.2424 30.08 / 0.9231 / 0.1960
Horse 25.26 / 0.9054 / 0.1272 27.53 / 0.9227 / 0.1108 25.70 / 0.9044 / 0.1321 25.85 / 0.9027 / 0.1311
Lighthouse 21.99 / 0.8482 / 0.2061 23.71 / 0.8701 / 0.1846 22.70 / 0.8556 / 0.2061 23.85 / 0.8627 / 0.1822
Mip-NeRF-360 Room 32.31 / 0.9401 / 0.1295 33.19 / 0.9507 / 0.1075 32.33 / 0.9370 / 0.1365 32.30 / 0.9454 / 0.1062
Kitchen 32.12 / 0.9370 / 0.1114 32.93 / 0.9429 / 0.1056 31.50 / 0.9303 / 0.1248 32.00 / 0.9365 / 0.1084
Bonsai 32.26 / 0.9495 / 0.0977 32.77 / 0.9554 / 0.0861 32.03 / 0.9460 / 0.1055 31.85 / 0.9464 / 0.0940
Counter 29.84 / 0.9210 / 0.1191 30.30 / 0.9303 / 0.1059 29.60 / 0.9142 / 0.1302 29.88 / 0.9201 / 0.1091
DTU Scan24 25.64 / 0.9399 / 0.0780 29.28 / 0.9531 / 0.0605 26.59 / 0.9402 / 0.0889 26.89 / 0.9332 / 0.0914
Scan37 25.81 / 0.9285 / 0.0973 27.58 / 0.9448 / 0.0757 25.81 / 0.9341 / 0.0870 24.73 / 0.9165 / 0.1019
Scan40 24.73 / 0.9278 / 0.1126 29.35 / 0.9528 / 0.0776 26.04 / 0.9218 / 0.1165 26.92 / 0.9323 / 0.0992
Scan55 28.48 / 0.9199 / 0.1750 30.66 / 0.9470 / 0.1093 28.59 / 0.9187 / 0.1834 29.50 / 0.9322 / 0.1331
Scan63 30.48 / 0.9677 / 0.0577 32.99 / 0.9767 / 0.0434 31.62 / 0.9671 / 0.0567 29.51 / 0.9528 / 0.0661
Scan65 27.69 / 0.9131 / 0.1913 30.65 / 0.9417 / 0.1332 29.30 / 0.9191 / 0.1982 29.59 / 0.9278 / 0.1476
Scan69 26.88 / 0.9142 / 0.1564 27.85 / 0.9306 / 0.1266 27.26 / 0.9154 / 0.1624 26.79 / 0.9168 / 0.1390
Scan83 29.64 / 0.9397 / 0.1876 30.59 / 0.9468 / 0.1670 29.69 / 0.9336 / 0.1975 30.08 / 0.9409 / 0.1831
Scan97 26.84 / 0.9180 / 0.1670 27.68 / 0.9246 / 0.1588 27.03 / 0.9172 / 0.1742 28.52 / 0.9272 / 0.1600
Scan105 30.23 / 0.9414 / 0.1511 32.26 / 0.9507 / 0.1381 30.06 / 0.9364 / 0.1659 31.23 / 0.9465 / 0.1479
Scan106 32.93 / 0.9443 / 0.1828 35.33 / 0.9628 / 0.1133 33.89 / 0.9469 / 0.1851 34.14 / 0.9502 / 0.1764
Scan110 31.67 / 0.9396 / 0.1945 33.49 / 0.9553 / 0.1530 31.88 / 0.9391 / 0.2011 32.52 / 0.9433 / 0.1812
Scan114 29.76 / 0.9266 / 0.1861 31.50 / 0.9518 / 0.1091 31.26 / 0.9366 / 0.1792 31.28 / 0.9383 / 0.1703
Scan118 33.40 / 0.9376 / 0.2074 36.00 / 0.9647 / 0.1142 34.40 / 0.9442 / 0.2063 35.37 / 0.9503 / 0.1813
Scan122 33.58 / 0.9436 / 0.2151 35.88 / 0.9626 / 0.1343 34.42 / 0.9435 / 0.2161 34.34 / 0.9487 / 0.1943
@@
Figure 7: Qualitative results. We provide rendering results of three more scenes from each dataset: Tanks&Temples [Lighthouse, Francis, Truck], Mip-NeRF-360 [Bansai, Kitchen, Counter], and DTU [Scan105, Scan118, Scan65]. For competitors, we use the maximum number of Gaussians recommended by the method. Our results are reported for 300K billboards.

### v4

@@ 4.3 Ablation study
In Table 5, we provide an ablation study of different aspects of our approach on Tanks&Temples dataset with $800\times 600$ resolution. The use of both alpha and RGB textures is essential for rendering quality. Excluding $T^{\alpha}$ significantly increases LPIPS, while excluding $T^{\textrm{RGB}}$ reduces PSNR and SSIM. Texture regularization term preserves over-fitting and leads to additional metrics improvement. It also encourages sparsity in textures and reduces storage costs in combination with compression. The use of compression leads to a slight metric reduction but allows to significantly reduce storage space requirements. Therefore, all parts of the proposed approach are essential to get the best result. However, disabling compression could further improve metrics.
@@ 7.1 Visual results
In Figure 7, we demonstrate additional renderings with ray-tracing effects and extracted meshes for the DTU dataset.
@@ 7.2 Detailed results
In Table 9, we report scene-by-scene results on the experiments provided in Table 1. We demonstrate state-of-the-art metrics quality for the Mip-NeRF-360 dataset on the indoor scenes while falling shortly behind for the outdoor scenes. This is caused by a huge number of small monochromic details (e.g. leafs) that are more efficient to represent with Gaussian primitives. For the Tanks&Temples outdoor scenes and DTU indoor scenes, we provide the best NVS for all cases.
7.3 Comparison to Texture-GS
[…]
Texture-GS 30.03 0.0135 0.1440
BBSplat(Ours) 30.81 0.0123 0.1055
7.4 Limitations and future work
One of the limitations of the proposed method is the NVS for the large outdoor scenes. While we provide high rendering quality for Tanks&Temples, there is still room for improvement in future works for the Mip-NeRF-360 dataset. Another limitation is training time, as it takes about 40 minutes to fit a scene, compared with 5 minutes for 3DGS. This slowdown is caused by backpropagation to the textures, and while it still significantly outperforms NeRF-based methods in terms of speed, it could be a bottleneck for some applications. Another limitation is that the use of a higher amount of primitives for outdoor scenes leads to over-fitting. In future works, we are going to focus on resolving these limitations.
@@ 7.4 Limitations and future work
Figure 7: Ray-tracing and mesh extraction. a) BBSplat representation of the scene as explicit textured planes allows the implementation of ray-tracing effects during rasterization (e.g. religtning or mirror effects). BBSplat representation can be rasterized as splatting or with classic rasterization tools (e.g. Blender). b) Rasterization of planes with alpha texture provide accurate objects suitable for mesh extraction. c) Alternatively, our representation can be used to extract accurate 3D meshes, surpassing the accuracy of 3D Gaussians-based methods (e.g. SuGaR).
@@
Tanks&Temples Train 19.81 / 0.7229 / 0.3347 21.21 / 0.7522 / 0.3120 19.35 / 0.7149 / 0.3417 21.38 / 0.7702 / 0.2604
Truck 20.43 / 0.7511 / 0.3192 23.91 / 0.8245 / 0.2513 20.77 / 0.7479 / 0.3283 24.19 / 0.8418 / 0.1957
Francis 24.16 / 0.8456 / 0.3368 27.82 / 0.8813 / 0.2964 22.61 / 0.8356 / 0.3485 28.56 / 0.9079 / 0.2334
Horse 19.17 / 0.7831 / 0.2973 24.93 / 0.8695 / 0.2080 19.92 / 0.7775 / 0.3116 24.72 / 0.8847 / 0.1604
Lighthouse 20.93 / 0.8044 / 0.2773 22.25 / 0.8163 / 0.2668 21.45 / 0.8053 / 0.2817 22.75 / 0.8332 / 0.2214
Mip-NeRF-360 Room 29.24 / 0.8924 / 0.2721 30.93 / 0.9056 / 0.2541 29.76 / 0.8914 / 0.2731 31.13 / 0.9146 / 0.2150
Kitchen 24.56 / 0.8712 / 0.2129 29.16 / 0.9003 / 0.1787 25.93 / 0.8745 / 0.2095 30.35 / 0.9116 / 0.1430
Bonsai 25.47 / 0.8856 / 0.2986 30.48 / 0.9230 / 0.2447 26.24 / 0.8895 / 0.2944 29.62 / 0.9085 / 0.2334
Counter 26.36 / 0.8716 / 0.2628 28.28 / 0.8910 / 0.2350 26.65 / 0.8701 / 0.2649 28.30 / 0.8941 / 0.2061
Bicycle 23.04 / 0.5912 / 0.4354 23.87 / 0.6675 / 0.3652 22.74 / 0.5738 / 0.4460 23.28 / 0.6520 / 0.3096
Garden 23.98 / 0.7093 / 0.3400 25.06 / 0.7512 / 0.2974 24.26 / 0.7053 / 0.3419 25.75 / 0.7873 / 0.1907
Stump 24.08 / 0.6387 / 0.4035 25.82 / 0.7201 / 0.3311 24.05 / 0.6399 / 0.4047 24.00 / 0.6547 / 0.3376
DTU Scan24 22.09 / 0.8450 / 0.2590 27.19 / 0.8745 / 0.2319 21.83 / 0.8423 / 0.2681 27.47 / 0.8850 / 0.1677
Scan37 20.01 / 0.8097 / 0.2942 24.18 / 0.8449 / 0.2673 20.83 / 0.8187 / 0.2881 24.93 / 0.8670 / 0.2127
Scan40 19.21 / 0.7676 / 0.3535 22.46 / 0.8132 / 0.3196 20.47 / 0.7717 / 0.3493 25.56 / 0.8709 / 0.2081
Scan55 25.25 / 0.7934 / 0.3831 27.88 / 0.8221 / 0.3775 25.67 / 0.7959 / 0.3857 27.74 / 0.8511 / 0.2746
Scan63 23.91 / 0.9098 / 0.2156 31.08 / 0.9411 / 0.1792 30.11 / 0.9228 / 0.1963 30.32 / 0.9422 / 0.1652
Scan65 27.76 / 0.8476 / 0.3648 29.41 / 0.8585 / 0.3605 29.32 / 0.8513 / 0.3688 29.67 / 0.8754 / 0.2803
Scan69 26.20 / 0.8485 / 0.3545 26.97 / 0.8592 / 0.3483 26.34 / 0.8465 / 0.3576 26.71 / 0.8660 / 0.2845
Scan83 28.36 / 0.8629 / 0.3888 29.62 / 0.8679 / 0.3818 28.69 / 0.8590 / 0.3921 28.97 / 0.8685 / 0.3604
Scan97 22.90 / 0.8263 / 0.3762 27.67 / 0.8588 / 0.3478 24.45 / 0.8360 / 0.3673 27.39 / 0.8579 / 0.3268
Scan105 24.58 / 0.8429 / 0.3781 30.10 / 0.8687 / 0.3579 23.92 / 0.8362 / 0.3797 29.35 / 0.8765 / 0.3154
Scan106 30.46 / 0.8780 / 0.3632 33.38 / 0.8983 / 0.3455 30.39 / 0.8723 / 0.3629 32.93 / 0.9020 / 0.3133
Scan110 30.71 / 0.8779 / 0.3773 31.55 / 0.8898 / 0.3666 31.12 / 0.8806 / 0.3771 31.42 / 0.8798 / 0.3324
Scan114 28.58 / 0.8642 / 0.3540 29.82 / 0.8807 / 0.3462 28.65 / 0.8636 / 0.3547 29.72 / 0.8870 / 0.2976
Scan118 31.77 / 0.8856 / 0.3561 34.06 / 0.9010 / 0.3482 32.49 / 0.8886 / 0.3572 34.41 / 0.9090 / 0.3077
Scan122 29.83 / 0.8640 / 0.3603 34.19 / 0.8949 / 0.3459 33.52 / 0.8956 / 0.3521 34.41 / 0.9069 / 0.3092
@@
Figure 8: Qualitative results. We provide rendering results of three more scenes from each dataset: Tanks&Temples [Lighthouse, Francis, Truck], Mip-NeRF-360 [Bansai, Kitchen, Counter], and DTU [Scan105, Scan118, Scan65]. For competitors, we use the maximum number of Gaussians recommended by the method.